OpenStudy (anonymous):

how can i proof g'(x)=1/F'(G(X)) if F(X) &G(X) are inverse of each other?

OpenStudy (psymon):

Well, in terms of a proof, how about this: If \[g(x)=f ^{-1}(x)\] Then let y = f^-1(x) Therefore f(y) = x and f'(y) = dx/dy. Since y = g(x), we can say f'(g(x)) = dx/dy. And if g(x) = y, then g'(x) = dy/dx, which is the same as: \[\frac{ 1 }{ (dx/dy) }=g'(x)\] If that makes sense o.o

OpenStudy (dumbcow):

haha im thinking how confusing that ^^ looks to a beginning calc student

OpenStudy (psymon):

I don't know the best way to like....dumb it down o.o

OpenStudy (psymon):

I would imagine the wiki page on it would be more confusing, though, lol.

OpenStudy (dumbcow):

me neither, it takes me a few minutes of just staring at it